# Training and Evaluating with Human Label Variation: An Empirical Study

### v1

@@ Abstract
Human label variation (HLV) challenges the standard assumption that an example has a single ground truth, instead embracing the natural variation in human labelling to train and evaluate models. While various training methods and metrics for HLV have been proposed, there has been no systematic meta-evaluation of HLV evaluation metrics, contributing to the lack of clarity in the best HLV training method. We propose new evaluation metrics and training methods and empirically meta-evaluate HLV evaluation metrics. We find that training on either disaggregated annotations or soft labels often performs best across metrics, and that our proposed soft metric correlates best with human preference. [^1] Code is available at: https://github.com/kmkurn/train-eval-hlv
@@ 1 Introduction
Human label variation (HLV) challenges the standard assumption that an example has a single ground truth (Plank, 2022). Without this assumption, model training and evaluation are no longer straightforward. To address this issue, recent work has proposed various methods to train models and metrics to evaluate their performance in the HLV context. However, there has not been a systematic meta-evaluation of HLV evaluation metrics, resulting in prior work reporting model performance across a wide variety of metrics, thwarting systematic model comparison. This has in turn contributed to a lack of clarity in the best HLV training method (Uma et al., 2021). In addition, NLP has largely ignored the field of remote sensing that has dealt with issues in model evaluation against a fuzzy ground truth.
In this work, we: (a) propose soft evaluation metrics, taking inspiration from remote sensing research; (b) propose new training methods based on the soft evaluation metrics; and (c) perform an empirical meta-evaluation of HLV evaluation metrics. Most existing metrics for HLV represent human judgements as probability distributions over classes and employ information-theoretic measures such as divergence. Taking inspiration from remote sensing research, we represent human judgement distributions as degrees of membership over fuzzy sets and generalise standard metrics such as accuracy into their soft versions using fuzzy set operations. To the best of our knowledge, we are the first to propose such soft metrics for HLV.
Because they are differentiable, we then use these soft metrics as training objectives. Experiments on 6 datasets show that training on either disaggregated annotations or soft labels often achieves the best performance across metrics. To perform the meta-evaluation, we recruit expert annotators and collect their preference on predicted judgement distributions. The meta-evaluation shows that our proposed soft metric correlates best with human preference, but the correlation is modest, highlighting the need for better evaluation metrics for HLV.
@@ 2 Related Work
First introduced by Plank (2022), the term human label variation (HLV) captures the fact that disagreeements in annotations can be well-founded and thus signal for data-driven methods. It challenges the traditional notion in machine learning that an instance has a single ground truth. Training methods that can handle such variation had been proposed before the term was coined (Sheng et al., 2008; Peterson et al., 2019; Uma et al., 2020, inter alia). Newer methods have also been employed in the Learning with Disagreement shared task (Leonardelli et al., 2023) which provides a benchmark for HLV. While a systematic investigation of these methods exists (Uma et al., 2020), it employs smaller pretrained language models and covers only binary or multiclass classification tasks. In contrast, our work additionally employs a large language model and covers multilabel tasks. More importantly, our work is the first to empirically meta-evaluate commonly-used evaluation metrics in the HLV context. In addition, we propose new soft metrics, which we also include in the meta-evaluation.
@@
While evaluating in the HLV context is a relatively new concept in NLP, evaluating against a fuzzy [^2] The terms “fuzzy” and “crisp” are sometimes used in remote sensing literature to mean “soft” and “hard” in NLP literature. reference is a well-studied area in remote sensing research. Early work by Foody (1996) used a measure of distance that is equivalent to twice the Jenson-Shannon divergence (Lin, 1991). Binaghi et al. (1999) argued that entropy-based measures are sensible only if the reference is crisp, and proposed the use of fuzzy set theory to compute the soft version of standard evaluation metrics such as accuracy. A similar approach was also proposed by Harju and Mesaros (2023) for audio data. Key to this approach is the minimum function used to compute the class membership scores given a fuzzy reference and a fuzzy model output. Other proposed functions include product (Lewis and Brown, 2001), sum (Pontius and Connors, 2006), and a composite of such functions designed to ensure diagonality of the confusion matrix when the reference and the model output match perfectly (Pontius and Cheuk, 2006; Silván-Cárdenas and Wang, 2008). Approaches that take the cost of misclassifications into account have also been proposed (Gómez et al., 2008). In this work, we take the fuzzy set approach by Binaghi et al. (1999) and apply it to text classification tasks.
@@ 3 Evaluation Metrics for HLV
Table 1 provides an overview of evaluation metrics used in this work.
@@ 3.1 Multiclass Metrics
    This addresses the issue with accuracy that it is biased toward the majority class when class proportions are imbalanced.
@@
    Novel to this work, for evaluation in the context of HLV. This is a special case of the soft micro F1 score proposed later in Section 3.2.
@@
    Also novel to this work, to address the issue of class imbalance with soft accuracy and other existing HLV metrics. [^3] In the context of HLV, we understand the concept of class imbalance analogously to that in the context where class assignments are hard: there exists a class $k$ such that $\sum_{i}P_{ik}\operatorname{>\!\!>}\sum_{i}P_{il}$ for all $l\neq k$. Note that this metric implies the existence of soft class-wise metrics (see Appendix A).
@@
    Jensen-Shannon divergence (Uma et al., 2021), which we modify into its positvely oriented version (PO-JSD for short) where:
@@
By noting that $N=\sum_{ik}P_{ik}$, soft accuracy can be interpreted as the proportion of the judgement distribution that is correctly predicted. This interpretation is similar to the hard counterpart (i.e. proportion of correctly predicted examples) and arguably more intuitive than that of existing metrics because it is stated directly in terms of the problem of predicting human judgement distribution. Other similarities include symmetry (due to the symmetry of $\min$) and boundedness (between 0 and 1), where the lower and the upper bounds are achieved if and only if $P_{ik}Q_{ik}=0$ and $P_{ik}=Q_{ik}$ respectively for all $i,k$. Note also that soft accuracy is reduced to the standard hard accuracy when all rows of both $\mathbf{P}$ and $\mathbf{Q}$ are one-hot vectors (i.e. no label variation).
@@
Other possible evaluation metrics include entropy similarity (Uma et al., 2021), (negative) cross-entropy (Peterson et al., 2019; Pavlick and Kwiatkowski, 2019), (negative) Kullback-Leibler divergence (Nie et al., 2020), (the positively oriented version of) information closeness (Foody, 1996) and (the positively oriented version of) Jensen-Shannon distance (Nie et al., 2020). However, they are excluded from this study because of their expected high correlation with either PO-JSD or entropy correlation.
@@ 3.2 Multilabel Metrics
where $J^{P}_{ik}=\mathds{1}(P_{ik}>0.5)$, and $J^{Q}_{ik}$ is defined analogously. Similarly, we define the soft version of the micro F1 score as follows (see derivations in Appendix A):
@@
3.3 Relationship between Metrics
To analyse the relationship between metrics, we compute Pearson correlation between each pair of multiclass metrics (excluding the macro ones) for various number of classes (see Appendix B for details), similar to prior work (Chicco and Jurman, 2020).
Soft Accuracy and PO-JSD
@@ 3.3 Relationship between Metrics
Despite their strong positive correlation, soft accuracy and PO-JSD have a notable difference: soft accuracy is upper-bounded by PO-JSD (see Appendix C for proof). This fact implies that soft accuracy penalises a model more heavily than PO-JSD, as illustrated in Figure 1. It shows that a skewed predicted judgement distribution of $(0.2,0.8)$ has a soft accuracy of $0.7$, much less than the PO-JSD ($>0.9$). The high PO-JSD is potentially misleading as it is very close to the maximum possible value of 1.
4 Training Methods for HLV
We experiment with a total of 14 HLV training methods, as summarised in Table 3. The first method (ReL) is the simplest and uses standard cross-entropy loss as the training objective based on prior work (Sheng et al., 2008). The next 8 methods (MV, AR, ARh, AL, SL, SLMV, AE, AEh) are from the Learning with Disagreements shared task (Leonardelli et al., 2023) which also use standard cross-entropy loss. The next 3 methods (JSD, SmF1, SMF1) use soft metrics that are differentiable as the training objective. The last 2 methods (LA-min, LA-max) aggregate the cross-entropy losses of all annotations for a given input.
@@ 4 Training Methods for HLV
Name Labels Language Annotators $N$ $J$ $K$ A M
HS-Brexit (Akhtar et al., 2021) Hate speech English Experts 1K $6.0$ 2
MD-Agreement (Leonardelli et al., 2021) Offensiveness English Crowds 10K $5.0$ 2
ArMIS (Almanea and Poesio, 2022) Misogyny Arabic Experts 1K $3.0$ 2
ChaosNLI (Nie et al., 2020) NLI English Crowds 3K $100.0$ 3
MFRC (Trager et al., 2022) Moral sentiment English Experts 18K $3.4$ 8
Table 4: Public datasets used in the empirical meta-evaluation of metrics, along with the number of examples to the nearest thousand ($N$), average number of annotations per example ($J$), number of classes ($K$), whether annotator identity is present (A), and whether the task is multilabel classification (M).
@@
    Training on the disaggregated labels directly (Sheng et al., 2008), which we refer to as repeated labelling (ReL). For example, if the training data only has a single instance $x$ with annotations $y_{1}$ and $y_{2}$ then ReL constructs new training data $\{(x,y_{1}),(x,y_{2})\}$.
@@
    Training on the majority-voted labels (MV). This is the baseline model given by the shared task organisers.
@@
    Instance weighting, where the weights are the sum of annotator ranking scores for that instance (Cui, 2023). Suppose that training instance $x_{1}$ is labelled as A, A, B, and A by 4 annotators; and $x_{2}$ is labelled as A, B, and B by only the first 3 annotators. There are two ways of computing the ranking score of an annotator:
@@
      Average judgement of the majority-voted label given by the annotator (AR). The annotator ranking scores are (in order) $\frac{3}{4}$, $\frac{1}{2}\left(\frac{3}{4}+\frac{2}{3}\right)$, $\frac{2}{3}$, and $\frac{3}{4}$.
@@
      Average number of times the annotator agrees with the majority (ARh). The scores are $\frac{1}{2}$, $\frac{2}{2}$, $\frac{1}{2}$, and $\frac{1}{1}$.
    For both methods, the ranking score is zero if the annotator never agrees with the majority. Majority-voted labels are used as ground truth for training. The training objective is the cross-entropy loss multiplied by the instance weights.
@@
    Ambiguous labelling (AL) for instances with disagreements (Gajewska, 2023). In short, if annotators are not unanimous in their judgement then the instance is labelled with all selected classes exactly once and included in the training set. Because it was developed for single-label tasks, we exclude this method from multilabel tasks. [^5] Extending this method to multilabel tasks is beyond the scope of this work.
@@
    Training with soft labels (SL) as targets (Maity et al., 2023; Wan and Badillo-Urquiola, 2023). For example, if an instance is labelled as positive by 80% of annotators then the log-likelihood is $0.8\log p(1)+0.2\log p(0)$, where $p(1)$ and $p(0)$ denote the predicted probability of the positive and the negative class respectively.
@@
    Multi-task learning using both the majority-voted and the soft labels as targets (SLMV; Grötzinger et al., 2023). Two models with a shared encoder are trained to predict the two types of labels respectively, which is similar to the method of Fornaciari et al. (2021). For example, if an instance is labelled as positive by 80% of annotators then the multitask log-likelihood is $0.8\log p_{\theta_{1}}(1)+0.2\log p_{\theta_{1}}(0)+\log p_{\theta_{2}}(1)$, where $\theta_{1}$ and $\theta_{2}$ denote the parameters of the first and the second model respectively. We predict only the soft labels for evaluation.
@@
    Modelling each annotator separately followed by ensembling (Sullivan et al., 2023; Vitsakis et al., 2023). In this approach, we have as many models as there are annotators. Each model is trained on the labels given by a distinct annotator. In practice, we share the encoder parameters across models. Suppose that there are 3 annotators, and the predicted judgements of the positive class for an instance are 60%, 30%, and 90%. At test time, there are 2 ways to ensemble these predictions:
@@
      Taking the average (AE; Sullivan et al., 2023). The final predicted judgement for the positive class is $\frac{1}{3}\left(0.6+0.3+0.9\right)$.
@@
      Computing the distribution of most likely labels (AEh; Vitsakis et al., 2023). The final predicted judgement for the positive class is $\frac{2}{3}$ because 2 out of 3 judgements are over 50%.
@@
    Training using JSD (JSD), $1-$ soft accuracy for multiclass or micro F1 for multilabel tasks (SmF1), and $1-$ soft macro F1 (SMF1) as the objective (see Section 3 for definitions). [^6] We exclude entropy correlation because it doesn’t have a unique maximiser. These approaches leverage the differentiability of soft metrics, resulting in a single objective for both training and inference. By directly optimising the evaluation metric at training time, we expect the models to exhibit superior performance, especially when evaluated with the corresponding metric.
@@
    Aggregating the losses of all annotations for a given training instance. For example, given a training instance $x$ with annotations $y_{1}$ and $y_{2}$, we compute the losses $l_{1}=\mathcal{L}(x,y_{1})$ and $l_{2}=\mathcal{L}(x,y_{2})$ where $\mathcal{L}$ is the cross-entropy loss function. Then, we update model parameters based on the gradient of $g(l_{1},l_{2})$ where $g$ is an aggregation function. We experiment with the $\min$ (LA-min) and the $\max$ (LA-max) functions as $g$. Using the $\min$ (resp. $\max$) function means selecting the least (resp. most) “surprising” annotation for the model. Using the mean aggregation is mathematically equivalent to SL and thus excluded.

### v2

@@ Abstract
Human label variation (HLV) challenges the standard assumption that a labelled instance has a single ground truth, instead embracing the natural variation in human annotation to train and evaluate models. While various training methods and metrics for HLV have been proposed, it is still unclear which methods and metrics perform best in what settings. We propose new evaluation metrics for HLV leveraging fuzzy set theory. Since these new proposed metrics are differentiable, we then in turn experiment with employing these metrics as training objectives. We conduct an extensive study over 6 HLV datasets testing 14 training methods and 6 evaluation metrics. We find that training on either disaggregated annotations or soft labels performs best across metrics, outperforming training using the proposed training objectives with differentiable metrics. We also show that our proposed soft metric is more interpretable and correlates best with human preference. [^1] Code is available at: https://github.com/kmkurn/train-eval-hlv
@@ 1 Introduction
Human label variation (HLV) challenges the standard assumption that an example has a single ground truth (Plank, 2022). Without this assumption, model training and evaluation are no longer straightforward. For example, standard metrics such as accuracy assume a single ground truth label exists for each instance. To address this issue, recent work (Pavlick and Kwiatkowski, 2019; Peterson et al., 2019; Nie, Zhou, and Bansal, 2020; Uma et al., 2020, 2021; Cui, 2023; Gajewska, 2023; Maity et al., 2023; Wan and Badillo-Urquiola, 2023, inter alia) has proposed various methods to train models and metrics to evaluate their performance in the HLV context. However, the lack of a systematic study means it is still unclear which methods and metrics perform best in what settings.
In natural language processing (NLP), most existing metrics for HLV represent human judgements as probability distributions over classes and employ information-theoretic measures such as divergence. It has largely ignored the field of remote sensing that has dealt with issues in model evaluation against fuzzy ground truths (Foody, 1996; Binaghi et al., 1999; Lewis and Brown, 2001; Pontius and Cheuk, 2006; Pontius and Connors, 2006; Silván-Cárdenas and Wang, 2008; Gómez, Biging, and Montero, 2008). In remote sensing, the task is to classify patches of a satellite image of a region into categories such as land, water, and so forth. As an image patch often contains multiple categories, the field has developed evaluation metrics that can handle this variation in the ground truth labels. Taking inspiration from remote sensing research, we represent human judgement distributions as degrees of membership over fuzzy sets and generalise standard metrics such as accuracy into their soft versions using fuzzy set operations. To the best of our knowledge, we are the first to propose employing soft metrics for HLV in NLP.
Because these soft metrics are differentiable, we next explore the use of these soft metrics as training objectives. Through an extensive study testing 14 training methods and 6 evaluation metrics on 6 HLV datasets, we show that training on either disaggregated annotations or soft labels often achieves the best performance across metrics. We also conduct a meta-evaluation on the evaluation metrics. Albeit modest, we find that our proposed soft metric correlates best with human preference. However, this suggests that there is still substantial future work to develop better evaluation metrics for HLV.
To summarise, we: (a) propose soft metrics to measure model performance in the context of HLV, taking inspiration from remote sensing research; (b) propose new training methods based on the soft evaluation metrics; and (c) conduct an extensive study covering 6 HLV datasets, 14 training methods and 6 evaluation metrics to understand the best training methods and evaluation metrics.
@@ 2 Related Work
First introduced by Plank (2022), the term human label variation (HLV) captures the fact that disagreeements in annotations can be well-founded and thus signal for data-driven methods. It challenges the traditional notion in machine learning that an instance has a single ground truth. Training methods that accommodate such variation had been proposed before the term was coined (Sheng, Provost, and Ipeirotis, 2008; Peterson et al., 2019; Uma et al., 2020, inter alia). Newer methods have also been proposed in the literature (Deng et al., 2023; Lee, An, and Thorne, 2023; Chen et al., 2024; Rodríguez-Barroso et al., 2024) and in the Learning with Disagreement shared task (Leonardelli et al., 2023) which provides a benchmark for HLV. While a systematic investigation of some of these methods exists (Uma et al., 2020), it employs smaller pretrained language models and covers only binary or multiclass classification tasks. In contrast, our work additionally employs a large language model and covers multilabel tasks. Furthermore, we also propose new soft evaluation metrics for HLV.
@@
While evaluating in the HLV context is a relatively new concept in NLP, evaluating against a fuzzy [^2] The terms “fuzzy” and “crisp” are sometimes used in remote sensing literature to mean “soft” and “hard” in NLP literature. reference is a well-studied area in remote sensing research. Early work by Foody (1996) used a measure of distance that is equivalent to twice the Jenson-Shannon divergence (Lin, 1991). Binaghi et al. (1999) argued that entropy-based measures are sensible only if the reference is crisp, and proposed the use of fuzzy set theory to compute the soft version of standard evaluation metrics such as accuracy. A similar approach was also proposed by Harju and Mesaros (2023) for audio data. Key to this approach is the minimum function used to compute the class membership scores given a fuzzy reference and a fuzzy model output. Other proposed functions include product (Lewis and Brown, 2001), sum (Pontius and Connors, 2006), and a composite of such functions designed to ensure diagonality of the confusion matrix when the reference and the model output match perfectly (Pontius and Cheuk, 2006; Silván-Cárdenas and Wang, 2008). Approaches that take the cost of misclassifications into account have also been proposed (Gómez, Biging, and Montero, 2008). In this work, we take the fuzzy set approach by Binaghi et al. (1999) and apply it to text classification tasks.
@@ 3 Evaluation Metrics for HLV
We now discuss the evaluation metrics relevant for HLV data, first on multiclass metrics (Section 3.1), then multilabel metrics (Section 3.2), and finally the relationship between these metrics (Section 3.3). In these discussions, we will also introduce our use of soft metrics and explain how they relate to their hard variant counterparts. Table 1 provides a summary of the existing and proposed/new evaluation metrics.
@@ 3.1 Multiclass Metrics
    This metric addresses the issue with accuracy that is biased toward the majority class when class proportions are imbalanced.
@@
    This metric is novel to this work, for evaluation in the context of HLV. This is a special case of the soft micro F1 score proposed later in Section 3.2.
@@
    This metric is also novel to this work, to address the issue of class imbalance with soft accuracy and other existing HLV metrics. [^3] In the context of HLV, we understand the concept of class imbalance analogously to that in the context where class assignments are hard: there exists a class $k$ such that $\sum_{i}P_{ik}\operatorname{>\!\!>}\sum_{i}P_{il}$ for all $l\neq k$. Note that this metric implies the existence of soft class-wise metrics (see Appendix).
@@
    Jensen-Shannon divergence (Uma et al., 2021), which we modify into its positively oriented version (PO-JSD for short) where:
@@
By noting that $N=\sum_{ik}P_{ik}$, soft accuracy can be interpreted as the proportion of the judgement distribution that is correctly predicted. This interpretation is similar to the hard counterpart (i.e. proportion of examples that are correctly predicted) and arguably more intuitive than that of existing metrics because it is stated directly in terms of the problem of predicting human judgement distribution. Other similarities include symmetry (due to the symmetry of $\min$) and boundedness (between 0 and 1), where the lower and the upper bounds are achieved if and only if $P_{ik}Q_{ik}=0$ and $P_{ik}=Q_{ik}$ respectively for all $i,k$. Note also that soft accuracy is reduced to the standard hard accuracy when all rows of both $\mathbf{P}$ and $\mathbf{Q}$ are one-hot vectors (i.e. no label variation).
@@
Other evaluation metrics we considered included entropy similarity (Uma et al., 2021), (negative) cross-entropy (Peterson et al., 2019; Pavlick and Kwiatkowski, 2019), (negative) Kullback-Leibler divergence (Nie, Zhou, and Bansal, 2020), (the positively oriented version of) information closeness (Foody, 1996) and (the positively oriented version of) Jensen-Shannon distance (Nie, Zhou, and Bansal, 2020). However, we ultimately excluded them from this study because of their expected high correlation with either PO-JSD or entropy correlation.
@@ 3.2 Multilabel Metrics
where $J^{P}_{ik}=\mathds{1}(P_{ik}>0.5)$, and $J^{Q}_{ik}$ is defined analogously. Similarly, we define the soft version of the micro F1 score as follows (see derivations in the Appendix):
@@
3.3 Relationship between Metrics
To analyse the relationship between metrics, we compute Pearson correlation between each pair of multiclass metrics (excluding the macro ones) similar to prior work (Chicco and Jurman, 2020). For a given number classes, we sample $B$ matrices $\mathbf{P}$ and $\mathbf{Q}$ and compute the value of each metric on these $B$ pairs of $\mathbf{P}$ and $\mathbf{Q}$. The Pearson correlations between pairs of metrics are then computed based on these $B$ data points. We set $B=500$ for computational reasons. We report the correlation coefficients in Table 2.
Soft Accuracy and PO-JSD
@@ 3.3 Relationship between Metrics
Despite their strong positive correlation, soft accuracy and PO-JSD have a notable difference: soft accuracy is upper-bounded by PO-JSD.
Theorem 3.1.
Proof 3.2.
$\sum_{k}\min(\pi a_{k},(1-\pi)b_{k})\leq\frac{1}{2}\left(H([\pi,1-\pi])-\operatorname{JSD}_{\pi}(\mathbf{a},\mathbf{b})\right)$
for all $i$. Doubling both sides and taking the average of this inequality over $i$ give the desired result.
This fact implies that soft accuracy penalises a model more heavily than PO-JSD, as illustrated in Figure 1. It shows that a skewed predicted judgement distribution of $(0.2,0.8)$ has a soft accuracy of $0.7$, much less than the PO-JSD ($>0.9$). The high PO-JSD is potentially misleading as it is very close to the maximum possible value of 1.
4 Training Methods for HLV
@@
We now discuss the training methods. In total, we experiment with 14 HLV training methods, as summarised in Table 3. The first method (ReL) is the simplest and uses standard cross-entropy loss as the training objective based on prior work (Sheng, Provost, and Ipeirotis, 2008). The next 8 methods (MV, AR, ARh, AL, SL, SLMV, AE, AEh) are from the Learning with Disagreements shared task (Leonardelli et al., 2023) which also use standard cross-entropy loss. The next 3 methods (JSD, SmF1, SMF1) use soft metrics that are differentiable as the training objective. The last 2 methods (LA-min, LA-max) aggregate the cross-entropy losses of all annotations for a given input. In detail:
@@
    ReL (repeated labelling) trains on the disaggregated labels directly (Sheng, Provost, and Ipeirotis, 2008). For example, if the training data only has a single instance $x$ with annotations $y_{1}$ and $y_{2}$ then ReL constructs new training data $\{(x,y_{1}),(x,y_{2})\}$.
@@
    MV trains on the majority-voted labels. This is the baseline model given by the shared task organisers.
@@
    AR (annotator ranking) and ARh (annotator ranking hard) weight each training instance equal to the sum of ranking scores of its annotators (Cui, 2023). Suppose that training instance $x_{1}$ is labelled as A, A, B, and A by 4 annotators; and $x_{2}$ is labelled as A, B, and B by only the first 3 annotators. Then, to compute the ranking score of an annotator:
@@
      AR computes the average judgement of the majority-voted label given by the annotator. The annotator ranking scores are (in order) $\frac{3}{4}$, $\frac{1}{2}\left(\frac{3}{4}+\frac{2}{3}\right)$, $\frac{2}{3}$, and $\frac{3}{4}$.
@@
      ARh computes the average number of times the annotator agrees with the majority. The scores are $\frac{1}{2}$, $\frac{2}{2}$, $\frac{1}{2}$, and $\frac{1}{1}$.
    For both methods, the ranking score is zero if the annotator never agrees with the majority. Given an instance, the training objective is maximising the probability of its majority-voted label(s), scaled by its weight as defined above.
@@
    AL (ambiguous labelling) labels ambiguous instances, i.e. instances where annotators are not unanimous in their judgement, with all selected classes exactly once (Gajewska, 2023). Each instance-label pair is included in the training set. For example, if training instance $x_{1}$ is labelled as $y_{1}$, $y_{1}$, and $y_{2}$ by 3 annotators, and $x_{2}$ is labelled as $y_{2}$, $y_{2}$, and $y_{2}$, then AL produces $\{(x_{1},y_{1}),(x_{1},y_{2}),(x_{2},y_{2})\}$ as the training data. Because it was developed for single-label tasks, we exclude this method from multilabel tasks. [^6] Extending this method to multilabel tasks is beyond the scope of this work.
@@
    SL (soft labelling) trains on soft labels as targets (Maity et al., 2023; Wan and Badillo-Urquiola, 2023). For example, if an instance is labelled as positive by 80% of annotators then the log-likelihood is $0.8\log p(1)+0.2\log p(0)$, where $p(1)$ and $p(0)$ denote the predicted probability of the positive and the negative class respectively.
@@
    SLMV performs multi-task learning using both the soft (SL) and the majority-voted (MV) labels as targets (Grötzinger, Heuschkel, and Drews, 2023). Two models with a shared encoder are trained to predict the two types of labels respectively, which is similar to the method of Fornaciari et al. (2021). For example, if an instance is labelled as positive by 80% of annotators then the multitask log-likelihood is $0.8\log p_{\theta_{1}}(1)+0.2\log p_{\theta_{1}}(0)+\log p_{\theta_{2}}(1)$, where $\theta_{1}$ and $\theta_{2}$ denote the parameters of the first and the second model respectively. We predict only the soft labels for evaluation.
@@
    AE (annotator ensembling) and AEh (annotator ensembling hard) model each annotator separately followed by ensembling (Sullivan, Yasin, and Jacobs, 2023; Vitsakis et al., 2023). In this approach, we have as many models as there are annotators. Each model is trained on the labels given by a distinct annotator. In practice, we share the encoder parameters across models. Suppose that there are 3 annotators, and the predicted judgements of the positive class for an instance are 60%, 30%, and 90%. To ensemble these predictions at test time:
@@
      AE computes a simple mean (Sullivan, Yasin, and Jacobs, 2023). The final predicted judgement for the positive class is $\frac{1}{3}\left(0.6+0.3+0.9\right)$.
@@
      AEh computes the distribution of most likely labels (Vitsakis et al., 2023). The final predicted judgement for the positive class is $\frac{2}{3}$ because 2 out of 3 judgements are over 50%.
@@
    JSD, SmF1 (soft micro F1), and SMF1 (soft macro F1) train using JSD, $1-$ soft accuracy for multiclass or soft micro F1 for multilabel tasks, and $1-$ soft macro F1 as the objective respectively (see Section 3 for definitions). [^7] We exclude entropy correlation because it doesn’t have a unique maximiser. These approaches leverage the differentiability of soft metrics, resulting in a single objective for both training and inference. By directly optimising the evaluation metric at training time, we expect the models to exhibit superior performance, especially when evaluated with the corresponding metric.
@@
    LA-min (loss aggregation $\min$) and LA-max (loss aggregation $\max$) aggregate the losses of all annotations for a given training instance using the $\min$ and the $\max$ functions respectively. For example, given a training instance $x$ with annotations $y_{1}$ and $y_{2}$, we compute the losses $l_{1}=\mathcal{L}(x,y_{1})$ and $l_{2}=\mathcal{L}(x,y_{2})$ where $\mathcal{L}$ is the cross-entropy loss function. Then, we update model parameters based on the gradient of $g(l_{1},l_{2})$ where $g$ is either the $\min$ or the $\max$ function. Using the $\min$ (resp. $\max$) function means selecting the least (resp. most) “surprising” annotation for the model. Using the mean aggregation is mathematically equivalent to SL and thus excluded.